if tan tita is 5/12 then the value of sin tita plus cos tita is
Answer by lwsshak3(11628)

if tan tita is 5/12 then the value of sin tita plus cos tita is
***
use x for tita
you are working with a 5-12-13 right triangle:
sin x=5/13
cos x=12/13
sin x + cos x=17/13